# DualLQR: Efficient Grasping of Oscillating Apples using Task Parameterized Learning from Demonstration

### v1

@@ Abstract
Learning from Demonstration offers great potential for robots to learn to perform agricultural tasks, specifically selective harvesting. One of the challenges is that the target fruit can be oscillating while approaching. Grasping oscillating targets has two requirements: 1) close tracking of the target during the final approach for damage-free grasping, and 2) the complete path should be as short as possible for improved efficiency. We propose a new method called DualLQR. In this method, we use a finite horizon Linear Quadratic Regulator (LQR) on a moving target, without the need of refitting the LQR. To make this possible, we use a dual LQR set-up, with an LQR running in two seperate reference frames. Through extensive simulation testing, it was found that the state-of-art method barely meets the required final accuracy without oscillations and drops below the required accuracy with an oscillating target. DualLQR was found to be able to meet the required final accuracy even with high oscillations, with an accuracy increase of 60% for high orientation oscillations. Further testing on a real-world apple grasping task showed that DualLQR was able to successfully grasp oscillating apples, with a success rate of 99%.
I Introduction
@@
Within agricultural tasks, LfD is valuable for selective harvesting. Selective harvesting is challenging to automate because of variation and specific detachment motions required for safe harvesting of different types of fruits [7]. In selective harvesting, manipulators are used in three phases: 1) approach to target fruit, 2) detachment of fruit from the plant, and 3) placement of fruit in a storage container [8]. In the approach phase, the manipulator must approach the fruit safely, without damaging the plant or the fruit. One of the challenges during the approach phase is a moving target fruit [9]. This moving can occur because of hitting a compliant branch during the harvest operation, because of removing other fruits from the same branch or because of wind. Due to the compliance in a branch, this can result in long-lasting oscillations of the fruit. To still be able to grasp the fruit, the robot needs to adapt to the oscillation. To deal with the oscillation of a target, visual servo control has been used, but has trouble dealing with occlusions and cases of large movement due to collisions between the end-effector and the fruit [10]. This challenge of oscillating targets has not yet been addressed and solved in harvesting robotics [9], let alone for LfD in selective harvesting.
@@
To combine these requirements, two reference frames should be used. The first frame will be placed at the pose of the target for close tracking during the final approach. The second frame will be placed at the initial pose of the end-effector for reduced tracking behaviour when further away. By transforming the demonstrations to each reference frame, the variation in the demonstrations in each reference frame can be used as information on when tracking is important. Task Parameterized Gaussian Mixture Models (TP-GMM) combine these requirements [16]. Task Parameterization (TP) allows a task to be learned using multiple reference frames, transforming the demonstrations to each reference frame. In each reference frame, a GMM can be used to learn the variability of the demonstrations. By using multiple Gaussians, the variability can be learned in higher detail, e.g. variation over time and space. To transform the TP-GMM to trajectories that can be used by a robot, a regression method needs to be used, e.g. Gaussian Mixture Regression (GMR) and/or Linear Quadratic Regulator (LQR) [17].
@@
Calinon et al. [16] proposed a TP-GMM and GMR, coupled to an infinite horizon LQR (InfLQR) to be able to deal with targets moving during execution. At each timestep, the method performs three steps: 1) combine the GMMs in each reference frame given the new poses of the reference frames, 2) perform GMR on the combined GMM to extract the end-effector goal and covariance matrix, and 3) calculate control commands based on the output of the GMR using the InfLQR. By using an InfLQR, the method can calculate control commands without calculating the remaining trajectory at each timestep. The authors showed that the algorithm could execute the task with a moving target. However, no analysis of the performance was performed. InfLQR was used to limit computation time. This was important since the moving target would result in changes in the combined GMM, which would mean that a finite horizon LQR would need to be refitted on the complete remaining part of the trajectory. However, a finite horizon LQR is preferred since it can consider the varying precision in the remaining part of the task [16].
@@
We propose a new method called DualLQR. In this method, we use a finite horizon LQR on a moving target, without the need of refitting the LQR, thereby reducing the computations in the real time loop. To make this possible, we use a dual LQR set-up, with an LQR running in two reference frames. On the GMM in each reference frame, GMR is performed to extract the mean and covariance matrix at each timestep. The LQR in that reference frame is then fitted to these means and covariance matrices. Since this is done in the reference frame, a moving target would only result in a changing current state from the view of the LQR. During execution, each LQR is used to calculate the desired control in that reference frame. These controls are combined using a weighted average based on the precision matrix of the GMR in that reference frame. For the apple harvesting task, one reference frame is the stationary initial pose of the end-effector and the other reference frame is the oscillating fruit. The effect of tight tracking during the final approach is achieved in two ways. First, each LQR will track tighter if the covariance is smaller, which will happen when the target pose is closer to the origin of the reference frame. Second, this effect is enhanced by using the weighted average of precision matrices. Therefore, the LQR in the reference frame with higher required precision will affect the combined control output more.
@@
To evaluate the proposed method, we performed comprehensive simulation experiments testing the proposed method DualLQR and the existing method InfLQR [16] and comparing their performance. In addition, we tested the proposed method on a real robot on a simplified apple grasping task. Our code is available https://github.com/WUR-ABE/DualLQR.
II Proposed method
The proposed method has a demonstration phase, a learning phase, and a reproduction phase [18]. In the demonstration phase, a set of movements was recorded and processed to be used as data for the learning phase. The movements were recorded using the end-effector pose. In the learning phase, the TP-GMM was fitted to demonstrations in each reference frame using expectation-maximization (EM) [18]. In the reproduction phase, we made several changes before and during the execution of a trajectory. In Section II-A, the steps before execution are described. In Section II-B, the steps during execution are described.
@@ II Proposed method
Figure 1: Steps before execution. In each step, light grey lines show the demonstrations. Each step was performed separately in the start and end frames. The start frame is indicated in blue and the end frame is indicated in orange. (a) shows each reference frame’s GMM, where dots indicate the means and ellipses the covariances of the Gaussian components. (b) shows each reference frame’s GMR, where the line indicates the mean of the path and ellipses show the covariance along the path. (c) shows each reference frame’s LQR. The $\hat{K}_{t}^{P}$ and $\hat{K}_{t}^{V}$ indicate the gains, which are tuned using the GMR and an additional control cost $\rho$.
@@
Figure 2: Processes in the control loop during execution visualized in 2D. (a) indicates the task in the world frame, shown in black. The blue frame indicates the start frame and the orange frame indicates the end frame, which oscillates, as the black arrow indicates. The green vector is the control output provided by DualLQR. (b) indicates the task in each reference frame $i$, transformed using ${}_{i}^{W}T_{t}$ and with the covariance matrix $\Sigma_{i,t}$ from the GMR. Because of this transformation, the oscillation of the end-frame results in an oscillation of the end-effector in the end-frame. (c) indicates each LQR, using the end-effector pose $x_{i,t}$ as input. (d) indicates the balancing between the two LQRs, using $u_{i,t}$ and $\Sigma_{i,t}$, resulting in control output $u_{t}$.
II-A Before execution
Fig. 1 shows the steps before execution. In Fig. 1, the black square corresponds to the start reference frame and the black triangle corresponds to the end reference frame. These poses were used to obtain the GMM in each reference frame, see Fig. 1a. Then, we performed a Gaussian Mixture Regression to extract the mean and covariance matrix at each timestep in each reference frame, shown in Fig. 1b. In our case, means and covariance matrices contained the end-effector pose. The number of timesteps were determined using the duration of the demonstrations and the desired frequency of the online control loop. Lastly, the LQR was fitted to the output of the GMR [17], shown in Fig. 1c. Here, the control cost of the LQR was set to optimize the behaviour [17]. A lower control cost results in tighter tracking of the target, but will also increases the distance travelled.
II-B During execution

### v3

@@ Abstract
Learning from Demonstration offers great potential for robots to learn to perform agricultural tasks, specifically selective harvesting. One of the challenges is that the target fruit can be oscillating while approaching. Grasping oscillating targets has two requirements: 1) close tracking of the target during the final approach for damage-free grasping, and 2) the complete path should be as short as possible for improved efficiency. We propose a new method called DualLQR. In this method, we use a finite horizon Linear Quadratic Regulator (LQR) on a moving target, without the need of refitting the LQR. To make this possible, we use a dual LQR set-up, with an LQR running in two separate reference frames. Through extensive simulation testing, it was found that the state-of-art method barely meets the required final accuracy without oscillations and drops below the required accuracy with an oscillating target. DualLQR, on the other hand, was found to be able to meet the required final accuracy even with high oscillations, while travelling the least distance. Further testing on a real-world apple grasping task showed that DualLQR was able to successfully grasp oscillating apples, with a success rate of 99%.
Keywords:
Learning from Demonstration Oscillating Apple grasping
1 Introduction
@@
Within agricultural tasks, LfD is valuable for selective harvesting. Selective harvesting is challenging to automate because of variation and specific detachment motions required for safe harvesting of different types of fruits [11]. In selective harvesting, manipulators are used in three phases: 1) approach to target fruit, 2) detachment of fruit from the plant, and 3) placement of fruit in a storage container [6]. In the approach phase, the manipulator must approach the fruit safely, without damaging the plant or the fruit. One of the challenges during the approach phase is a moving target fruit [1]. This moving can occur because of hitting a compliant branch during the harvest operation, because of removing other fruits from the same branch, or because of wind. Due to the compliance in a branch, this can result in long-lasting oscillations of the fruit. To still be able to grasp the fruit, the robot needs to adapt to the oscillation. To deal with the oscillation of a target, visual servo control has been used, but has trouble dealing with unforeseen interactions [18]. This challenge of oscillating targets has not yet been addressed and solved in harvesting robotics [1], let alone for LfD in selective harvesting.
@@
To combine these requirements, the task can be learned relative to the pose of the moving target for close tracking during the final approach, and relative to the initial pose of the end-effector for reduced tracking behaviour when further away. These poses are then used as reference frames to transform the demonstrations to. By transforming the demonstrations to each reference frame, the variation in the demonstrations in each reference frame can be used as information on when tracking is important. Task Parameterized Gaussian Mixture Models (TP-GMM) combine these requirements [4]. Task Parameterization (TP) allows a task to be learned using multiple reference frames, transforming the demonstrations to each reference frame. In each reference frame, a GMM can be used to learn the variability of the demonstrations. By using multiple Gaussians, the variability can be learned in higher detail, e.g. variation over time and space. To transform the TP-GMM to trajectories that can be used by a robot, a regression method needs to be used, e.g. Gaussian Mixture Regression (GMR) and/or Linear Quadratic Regulator (LQR) [5].
@@
Calinon et al. [4] proposed a TP-GMM and GMR, coupled to an infinite horizon LQR (InfLQR) to deal with targets moving during execution. At each timestep, the method performs three steps: 1) combine the GMMs in each reference frame given the new poses of the reference frames, 2) perform GMR on the combined GMM to extract the end-effector goal and covariance matrix, and 3) calculate control commands based on the output of the GMR using the InfLQR. By using an InfLQR, the method can calculate control commands without calculating the remaining trajectory at each timestep. The authors showed that the algorithm could execute the task with a moving target. However, no analysis of the performance was performed. InfLQR was used to limit computation time. This was important since the moving target would result in changes in the combined GMM, which would mean that a finite horizon LQR would need to be refitted on the complete remaining part of the trajectory. However, a finite horizon LQR is preferred since it can consider the varying precision in the remaining part of the task [4].
@@
We propose a new method called DualLQR [^1] Our code is available at https://github.com/WUR-ABE/DualLQR. In this method, we use a finite horizon LQR on a moving target, without the need of refitting the LQR, thereby reducing the computations in the real time loop. To enable this, we use a dual LQR set-up, with an LQR running in two reference frames. During execution, each LQR is used to calculate the desired control in that reference frame. These controls are combined using a weighted average based on the precision matrix of the GMR in that reference frame. For the apple harvesting task, one reference frame is the stationary initial pose of the end-effector and the other reference frame is the oscillating fruit. The effect of tight tracking during the final approach is achieved in two ways. First, each LQR will track tighter if the covariance is smaller, which will happen when the target pose is closer to the origin of the reference frame. Second, this effect is enhanced by using the weighted average of precision matrices. Therefore, the LQR in the reference frame with higher required precision will affect the combined control output more.
@@
To evaluate the proposed method, we performed comprehensive simulation experiments testing the proposed method DualLQR and two existing methods, InfLQR [4] and using the LQR using a single reference frame. In addition, we tested the proposed method on a real robot on a simplified apple grasping task.
2 Background
In this section, we explain the existing methods on which DualLQR builds. First, we describe the Task-Parameterization to describe the task in reference frames. Next, we describe the InfLQR method [4]. Finally, we describe the LQR using a single reference frame.
2.1 Task-Parameterization
In this section, we explain the Task-Parameterization (TP). Initially, demonstrations were recorded, resulting in trajectories $X_{g}=[x_{1},\dots,x_{T}]\in\mathbb{R}^{D\times T}$ in a global frame $g$ and consisted of $T$ timesteps. In our case, $D=6$ consisting of a 3-dimensional Cartesian position and 3-dimensional Tait-Bryan angles for each timestep. These demonstrations were converted into reference frame $j$ using transformation matrices, resulting in trajectories $X_{j}$. In the reference frame $j$, the task can be encoded as $\xi_{j}=[\tau,X_{j}]^{\top}$, with $\tau=[1,\dots,T]$ and $\xi_{j}\in\mathbb{R}^{\left(1+D\right)\times T}$.
The task can be learned in all reference frames $J$ using a single GMM by combining the task encodings of each reference frame into the encoding $\xi=[\tau,X_{1},\dots,X_{J}]^{\top}$, with $\xi\in\mathbb{R}^{\left(1+J\cdot D\right)\times T}$
In order to combine reference frames, the GMMs need to be transformed into the same reference frame. To transform a GMM from reference frame $j$ into the global frame $g$, a linear transformation is performed. Each reference frames can be defined as an origin vector $b\in\mathbb{R}^{D+1}$ and a transformation matrix $A\in\mathbb{R}^{\left(D+1\right)\times\left(D+1\right)}$. For reference frame $j$, this is defined as [4]:
$\xi_{j}=\begin{bmatrix}\tau\\ X^{p}_{j}\\ X^{r}_{j}\end{bmatrix}\text{, }b_{j}=\begin{bmatrix}0\\ x^{p}_{j}\\ x^{r}_{j}\end{bmatrix}\text{, }A_{j}=\begin{bmatrix}I&0&0\\ 0&R_{j}&0\\ 0&0&I\end{bmatrix}$ (1)
where $\tau$ is the time dimension, $X^{p}_{j}$ is the trajectory of the 3-dimensional Cartesian position and $X^{r}_{j}$ is the trajectory of the 3-dimensional orientation represented using Tait-Bryan angles in frame $j$. $x^{p}_{j}$ indicates the 3-dimensional position and $x^{r}_{j}$ and $R_{j}$ indicate the 3-dimensional orientation, using the Tait-Bryan angles and rotation matrix of frame $j$. 00 are zeros vectors/matrices of the corresponding size and $I$ identity matrices of the corresponding size.
2.2 InfLQR
In this section, we describe how the infinite horizon LQR is initialized and fitted before execution, and how the method is used during execution.
2.2.1 Before execution
In InfLQR, the GMM is fitted before execution. A GMM combines $K$ Gaussian distributions in order to represent the demonstrated trajectories. The component $C_{k}$ is a Gaussian distribution $\mathcal{N}(\mu_{k},\Sigma_{k})$, using the task encoding $\xi$. These Gaussian components were fitted to the data using an Expectation-Maximization (EM) algorithm [3].
2.2.2 During execution
At each timestep $t$ during the execution, several steps need to be performed to get a GMM in the global frame. As the task was encoded using $\xi$, the GMM containing multiple references frames needs to be split into separate GMMs containing a single reference frames, with a task encoding of $\xi_{j}$ and $K$ components consisting of $\mathcal{N}(\mu_{k,j},\Sigma_{k,j})$.
As the task is still encoded in each reference frame, the next step is to obtain the global encoding given the current timestep $t$ and using the current location of the reference frames $A_{t,j}$ and $b_{t,j}$, defined as $\xi_{t}=[\tau,X_{g}]$. This is calculated using the product of the linearly transformed GMM from each reference frame $J$, using the following equations [4]:
$\Sigma_{k,t}=\left(\sum_{j=1}^{J}\left(A_{t,j}\Sigma_{k,j}A_{t,j}^{\top}\right)^{-1}\right)^{-1}$ (2)
$\mu_{k,t}=\Sigma_{k,t}\left(\sum_{j=1}^{J}\left(A_{t,j}\Sigma_{k,j}A_{t,j}^{\top}\right)^{-1}\left(A_{t,j}\mu_{k,j}+b_{t,j}\right)\right)$ (3)
where $\mu_{k,j}$ and $\Sigma_{k,j}$ are first transformed into the global reference using $A_{t,j}$ and $b_{t,j}$, after which the reference frames are added together based on the required precision in each frame. This results in a combined GMM for timestep $t$ with task encoding of $\xi_{t}$ and $K$ components consisting of $\mathcal{N}(\mu_{k,t},\Sigma_{k,t})$.
The next step is to calculate the target mean and covariance given the combined GMM. For this step, Gaussian Mixture Regression (GMR) is used. GMR uses the components of the GMM to determine the mean and covariance of the output dimensions $\mathcal{O}$ given a value in the input dimension $\mathcal{I}$. In this case, the input dimension corresponds to the time dimensions in the task encoding and the output dimension corresponds to the pose in the task encoding. The input and output dimensions are also applied to the components of the GMM, resulting in the following encoding, mean and covariance:
$\xi_{t}=\begin{bmatrix}\xi_{t}^{\mathcal{I}}\\ \xi_{t}^{\mathcal{O}}\end{bmatrix}\text{, }\mu_{k,t}=\begin{bmatrix}\mu_{k,t}^{\mathcal{I}}\\ \mu_{k,t}^{\mathcal{O}}\end{bmatrix}\text{, }\Sigma_{k,t}=\begin{bmatrix}\Sigma_{k,t}^{\mathcal{I}}&\Sigma_{k,t}^{\mathcal{IO}}\\ \Sigma_{k,t}^{\mathcal{OI}}&\Sigma_{k,t}^{\mathcal{O}}\end{bmatrix}$ (4)
At timestep $t$, $\xi_{t}^{\mathcal{I}}=t$, and can be used to calculate the target mean and covariance of the pose, using the following equations [4]:
$\hat{\mu}_{t}^{\mathcal{O}}=\sum_{k=1}^{K}h_{k}\left(\xi_{t}^{\mathcal{I}}\right)\left(\mu_{k,t}^{\mathcal{O}}+\Sigma_{k,t}^{\mathcal{OI}}{\Sigma_{k,t}^{\mathcal{I}}}^{-1}\left(\xi_{t}^{\mathcal{I}}-\mu_{k,t}^{\mathcal{I}}\right)\right)$ (5)
$\hat{\Sigma}_{t}^{\mathcal{O}}=\sum_{k=1}^{K}h_{k}\left({\xi_{t}^{\mathcal{I}}}\right)\left(\Sigma_{k,t}^{\mathcal{O}}-\Sigma_{k,t}^{\mathcal{OI}}{\Sigma_{k,t}^{\mathcal{I}}}^{-1}\Sigma_{k,t}^{\mathcal{IO}}\right)$ (6)
which depend on the activation function $h_{k}$:
$h_{k}\left(\xi_{t}^{\mathcal{I}}\right)=\frac{\pi_{k}\mathcal{N}\left(\xi_{t}^{\mathcal{I}}|\mu_{k,t}^{\mathcal{I}},\Sigma_{k,t}^{\mathcal{I}}\right)}{\sum_{k=1}^{K}\pi_{k}\mathcal{N}\left(\xi_{t}^{\mathcal{I}}|\mu_{k,t}^{\mathcal{I}},\Sigma_{k,t}^{\mathcal{I}}\right)}$ (7)
which indicates the importance of each component $k$ of the combined GMM given the input $\xi_{t}^{\mathcal{I}}$.
Next, the infinite horizon LQR is used to determine the required control outputs to achieve the target mean. The system in the LQR is defined as [5]:
$x_{t+1}=Ax_{t}+Bu_{t}$ (8)
In our case, $A=I$ and $B=I\cdot\Delta t$, with $\Delta t$ the size of the timestep. The infinite horizon LQR balances the control outputs $u_{t}$ using the matrices $R$ for the cost and $Q_{t}$ for the required precision [3]:
$R=I\cdot 10^{\rho}\text{, }Q_{t}=\left(\hat{\Sigma}_{t}^{\mathcal{O}}\right)^{-1}$ (9)
where $R\in\mathbb{R}^{D\times D}$ is constant and can be tuned using control cost $\rho$ and $Q_{t}$ is the precision matrix of the GMM. The control $u_{t}$ is then determined using the following equation [4]:
$u_{t}=K_{t}^{\mathcal{P}}\left(\hat{\mu}_{t}^{\mathcal{O}}-x_{t}\right)$ (10)
where $\hat{\mu}_{t}^{\mathcal{O}}$ is estimated by the GMR and $x_{t}$ indicates the state of the robot. The feedback gain $K_{t}^{\mathcal{P}}$ is calculated using the following equation [5]:
$K_{t}^{\mathcal{P}}=\left(R+B^{\top}P_{t}B\right)^{-1}B^{\top}P_{t}A$ (11)
which depends on the outcome of the Riccati equation:
$P_{t}=Q_{t}-A^{\top}\left(Q_{t}B\left(B^{\top}Q_{t}B+R\right)^{-1}B^{\top}Q_{t}-Q_{t}\right)A$ (12)
which for the infinite horizon LQR only considers the required precision at time $t$. With these equations, the feedback gain $K_{t}^{\mathcal{P}}$ will be time-variable and perform tighter tracking of target mean $\hat{\mu}_{t}^{\mathcal{O}}$ based on the desired precision $Q_{t}$.
2.3 Single LQR
Alternatively to combining the reference frames, the task could be performed in a single reference frame using a finite horizon LQR coupled to the moving target. This finite horizon LQR can then be fitted before execution.
2.3.1 Before execution
With a single LQR, the task is encoded as $\xi_{j,t}=[t,x_{j,t}^{\top}]^{\top}$. The GMM was fitted using EM. Since the task is consisted in the frame of the moving target, the GMR can be performed before execution and a finite horizon LQR can be fitted to this trajectory.
For the GMR, Equations 5 and 6 are used to calculate the desired trajectory $\hat{\mu}^{\mathcal{O}}=[\hat{\mu}_{0}^{\mathcal{O}}\cdots\hat{\mu}_{T}^{\mathcal{O}}]$, with variation $\hat{\Sigma}^{\mathcal{O}}=[\hat{\Sigma}_{0}^{\mathcal{O}}\cdots\hat{\Sigma}_{T}^{\mathcal{O}}]$.
The LQR is then fitted to this trajectory to determine the optimal control outputs to achieve the final target. The finite horizon LQR uses a set of recursive functions to determine the feedforward gain $K_{t}^{\mathcal{V}}$ and feedback gain $K_{t}^{\mathcal{P}}$ [5]:
$K_{t}^{\mathcal{V}}=\left(R+B^{\top}S_{t+1}B\right)^{-1}B^{\top}$ (13)
$K_{t}^{\mathcal{P}}=\left(R+B^{\top}S_{t+1}B\right)^{-1}B^{\top}S_{t+1}A$ (14)
$S_{t}=A^{\top}v_{t+1}\left(A-BK_{t}^{\mathcal{P}}\right)+Q_{t}$ (15)
$v_{t}=\left(A-BK_{t}^{\mathcal{P}}\right)^{\top}v_{t+1}+Q_{t}\hat{\mu}_{t}^{\mathcal{O}}$ (16)
where $A$ and $B$ describe the system as defined in Equation 8, $R$ is the control cost and $Q$ is the required precision, both defined in Equation 9. $S_{t}$ and $v_{t}$ are the Riccati and output function respectively, which are used to recursively calculate the gains from the final conditions $S_{T}=Q_{T}$ and $v_{T}=Q_{T}\hat{\mu}_{T}^{\mathcal{O}}$.
With these equations, the final conditions $S_{T}$ and $v_{T}$ are used to determine intermediate conditions and gains $K_{t}^{\mathcal{V}}$ and $K_{t}^{\mathcal{P}}$. These are used by the LQR during the execution to determine control outputs which achieve both the current conditions, but also take the final condition into account.
2.3.2 During execution
At each timestep, the control output can be directly determined from the gains and the output function, calculated in Equations 13, 14, and 16. In addition, the current state $x_{t}$ is used. This is the affine optimal control equation [5]:
$u_{i}=-K_{t}^{\mathcal{P}}x_{t}+K_{t}^{\mathcal{V}}v_{t+1}$ (17)
where the feedback term $K_{t}^{\mathcal{P}}$ allows the control to be determined in terms of the current state $x_{t}$ and the feedforward term $K_{t}^{\mathcal{V}}$ anticipates the desired reference signal.
3 Proposed method
When the task is only performed in the frame of the moving target, the number of performed equations during execution is reduced a lot. However, because the task is only performed relative to the target, the controller will always attempt to track the target, which can only be minimized by increasing the cost for tracking. On the other hand, the infinite horizon LQR can take both reference frames into account and thereby reduce the initial amount of tracking, but requires several complex computations for this.
Therefore, we propose DualLQR, which maintains both reference frames, while only adding minimal additional computation costs compare to using a single reference frame.
@@ 3 Proposed method
(a) Steps before execution. In each step, light grey lines show the demonstrations. Each step was performed separately in the start and end frames. The start frame is indicated in blue and the end frame is indicated in orange. (1) shows each reference frame’s GMM, where dots indicate the means and ellipses the covariances of the Gaussian components. (2) shows each reference frame’s GMR, where the line indicates the mean of the path and ellipses show the covariance along the path. (3) shows each reference frame’s LQR. The $K_{t}^{\mathcal{P}}$ and $K_{t}^{\mathcal{V}}$ indicate the gains, which are tuned using the GMR and an additional control cost $\rho$.
@@
(b) Processes in the control loop during execution visualized in 2D. (1) indicates the task in the world frame, shown in black. The blue frame indicates the start frame and the orange frame indicates the end frame, which oscillates, as the black arrow indicates. The green vector is the control output provided by DualLQR. (2) indicates the task in each reference frame $j$, transformed using ${}_{j}^{g}T_{t}$ and with the covariance matrix $\hat{\Sigma}_{j,t}^{\mathcal{O}}$ from the GMR. Because of this transformation, the oscillation of the end-frame results in an oscillation of the end-effector in the end-frame. (3) indicates each LQR, using the end-effector pose $x_{j,t}$ as input. (4) indicates the balancing between the two LQRs, using $u_{j,t}$ and $\hat{\Sigma}_{j,t}^{\mathcal{O}}$, resulting in control output $u_{t}$.
Figure 1: Flowcharts of the pre-process and real-time processes of DualLQR
3.0.1 Before execution
For DualLQR, the GMM encoded the task as $\xi_{t}=[t,x_{1,t}^{\top},x_{2,t}^{\top}]^{\top}$. The GMM is fitted using EM. Next, a separate GMM is created for each reference frame by splitting along the task dimensions, resulting $\xi_{1,t}=[t,x_{1,t}^{\top}]^{\top}$ and $\xi_{2,t}=[t,x_{2,t}^{\top}]^{\top}$. Next, the GMR and LQR are fitted in each reference frame, as described in section 2.3.1. The steps performed before execution are shown in Fig. 1a.
3.0.2 During execution